OpenStudy (anonymous):

find the next three terms in the sequence. 30, 22, 14, 6, . . . -3, -12, -21 -1, -8, -15 -1, -2, 3 -2, -10, -18

OpenStudy (warriorz13):

D

OpenStudy (anonymous):

thanks

OpenStudy (warriorz13):

30-8=22-8=14-8=6-=-2

OpenStudy (warriorz13):

NP